OpenStudy (anonymous):

Help Trigonometric substitution \[\int\limits_{0}^{5/2}\frac{ dx }{ \sqrt{25-x ^{2}} }\]

hartnn (hartnn):

so which substitution u tried ? hint : for a^2-x^2 try x = a sin t

OpenStudy (anonymous):

ok so \[\frac{ dx }{ \sqrt{25-(5\sin \theta)^{2} }}\]

hartnn (hartnn):

you need to find dx too!

hartnn (hartnn):

x = 5 sin t dx = .... ?

OpenStudy (anonymous):

dx=5cost

OpenStudy (anonymous):

so \[\frac{ 5\cos \theta }{ \sqrt{25\cos ^{2}\theta } }\] correct?

hartnn (hartnn):

yes!

OpenStudy (anonymous):

and now what?

hartnn (hartnn):

simplify the denominator! whats square root of 25 ?

OpenStudy (anonymous):

so 5cos t in the denominator?

hartnn (hartnn):

yes!

OpenStudy (anonymous):

so wouldn't that simplify to 1?

hartnn (hartnn):

absolutely.

hartnn (hartnn):

but we r not done yet!

hartnn (hartnn):

we need to change the limits of integration too, right ???

OpenStudy (anonymous):

to?

OpenStudy (anonymous):

pi/2, -pi/2?

hartnn (hartnn):

how ? when x= 0 what is t ? x = 5 sin t

OpenStudy (anonymous):

0

hartnn (hartnn):

correct! when x= 5/2 what is t ? x = 5 sin t

OpenStudy (anonymous):

pi/6

OpenStudy (anonymous):

and there's the answer

hartnn (hartnn):

yes, as integral of 1 is just x and then u put the limits pi/6-0 = pi/6

OpenStudy (anonymous):

oh I see I totally forgot that this was an integral question at its heart. Besides that my substitution skills were a little lacking that is. Ok this makes sense now

OpenStudy (anonymous):

Thanks for always being so helpful.

hartnn (hartnn):

you're welcome ^_^